### v1

# Matrix invertible extensions over commutative rings. Part I: general theory

Abstract.
A unimodular $2\times 2$ matrix with entries in a commutative $R$ is called extendable (resp. simply extendable) if it extends to an invertible $3\times 3$ matrix (resp. invertible $3\times 3$ matrix whose $(3,3)$ entry is 00). We obtain necessary and sufficient conditions for a unimodular $2\times 2$ matrix to be extendable (resp. simply extendable) and use them to study the class $E_{2}$ (resp. $SE_{2}$) of rings $R$ with the property that all unimodular $2\times 2$ matrices with entries in $R$ are extendable (resp. simply extendable). We also study the larger class $\Pi_{2}$ of rings $R$ with the property that all unimodular $2\times 2$ matrices of determinant 00 and with entries in $R$ are (simply) extendable (e.g., rings with trivial Picard groups or pre-Schreier domains). Among Dedekind domains, polynomial rings over $\mathbb{Z}$ and Hermite rings, only the EDRs belong to the class $E_{2}$ or $SE_{2}$. If $as(R)\leq 2$, then $R$ is an $E_{2}$ ring iff it is an $SE_{2}$ ring.
Key words and phrases:
ring, matrix, projective module, stable range, unimodular.
2020 Mathematics Subject Classification:
Primary: 15A83, 13G05, 19B10. Secondary: 13A05, 13F05, 15B33.
1. Introduction
Let $R$ be a commutative ring with identity; we denote by $U(R)$ its group of units, by $J(R)$ its Jacobson radical, and by $Pic(R)$ its Picard group. For $m,n\in\mathbb{N}=\{1,2,\ldots\}$, let $\mathbb{M}_{m\times n}(R)$ be the $R$-module of $m\times n$ matrices with entries in $R$; we view $\mathbb{M}_{n}(R):=\mathbb{M}_{n\times n}(R)$ as an $R$-algebra with identity $I_{n}$. Let ${GL}_{n}(R)$ be the general linear group of units of $\mathbb{M}_{n}(R)$, and let ${SL}_{n}(R):=\{M\in GL_{n}(R)|\det(M)=1\}$ be the special linear subgroup of ${GL}_{n}(R)$.
For a free $R$-module $F$, let $Um(F)$ be the set of unimodular elements of $F$, i.e., of elements $v\in F$ for which there exists an $R$-linear map $L:F\rightarrow R$ such that $L(v)=1$; we have an identity $U(R)=Um(R)$ of sets.
For $A\in\mathbb{M}_{n}(R)$, let $A^{T}$ be its transpose, let $\textup{Tr}(A)$ be its trace, let $\chi_{A}(\lambda)\in R[\lambda]$ be its (monic) characteristic polynomial, and let $\nu_{A}:=\chi_{A}^{\prime}(0)\in R$. E.g., if $n=3$, then $\chi_{A}(\lambda)=\lambda^{3}-\textup{Tr}(A)\lambda^{2}+\nu_{A}\lambda-\det(A)$.
As we will be using pairs, triples and quadruples extensively, the elements of $R^{n}$ will be denoted as $n$-tuples except for the case of the $R$-linear map $L_{A}:R^{n}\rightarrow R^{n}$ defined by $A$, in which case they will be viewed as $n\times 1$ columns. Let $\textup{Ker}_{A}$, $\textup{Coker}_{A}$ and $\textup{Im}_{A}$ be the kernel, cokernel and image (respectively) of $L_{A}$.
Recall that $B,C\in\mathbb{M}_{m\times n}(R)$ are said to be equivalent if there exist matrices $M\in{GL}_{m}(R)$ and $N\in GL_{n}(R)$ such that $C=MBN$; if $m=n$ and we can choose $N=M^{-1}$, then $B$ and $C$ are said to be similar. If all entries of $B-C$ belong to an ideal $I$ of $R$, then we say that $B$ and $C$ are congruent modulo $I$. By the reduction of $A\in\mathbb{M}_{n}(R)$ modulo $I$ we mean the image of $A$ in $\mathbb{M}_{n}(R/I)\cong\mathbb{M}_{n}(R)/\mathbb{M}_{n}(I)$.
We say that $B\in\mathbb{M}_{m\times n}(R)$ admits diagonal reduction if it is equivalent to a matrix whose off diagonal entries are 00 and whose diagonal entries $b_{1,1},\ldots,b_{s,s}$, with $s:=\min\{m,n\}$, are such that $b_{i,i}$ divides $b_{i+1,i+1}$ for all $i\in\{1,\ldots,s-1\}$.
We will use the shorthand ‘iff’ for ‘if and only if’ in all that follows.
For $Q\in\mathbb{M}_{3}(R)$, let $\Theta(Q)\in\mathbb{M}_{2}(R)$ be obtained from $Q$ by removing the third row and the third column. If $Q\in GL_{3}(R)$, then $\Theta(Q)$ modulo each maximal ideal of $R$ is nonzero, hence $\Theta(Q)\in Um(\mathbb{M}_{2}(R))$. The rule $Q\rightarrow\Theta(Q)$ defines a map
$\Theta:=\Theta_{R}:{SL}_{3}(R)\rightarrow Um(\mathbb{M}_{2}(R))$
and in this paper we study the image of $\Theta$ and in particular the class of rings $R$ for which $\Theta$ is surjective (i.e., it has a right inverse).
If $M=\left[\begin{array}[]{cc}a&b\\ c&d\end{array}\right]\in{GL}_{2}(R)$, let $\sigma(M):=\left[\begin{array}[]{ccc}a&b&0\\ c&d&0\\ 0&0&\det(M)^{-1}\end{array}\right]\in{SL}_{3}(R)$. Clearly $\Theta(\sigma(M))=M$, hence $GL_{2}(R)\subseteq{Im}(\Theta)$.
Definition 1.1.
We say that a matrix $A\in\mathbb{M}_{2}(R)$ is $SL_{3}$-extendable if there exists $A^{+}\in{SL}_{3}(R)$ such that $A=\Theta(A^{+})$, and we call $A^{+}$ an $SL_{3}$-extension of $A$. If we can choose $A^{+}$ such that its $(3,3)$ is 00, then we say that $A$ is simply $SL_{3}$-extendable and that $A^{+}$ is a simple $SL_{3}$-extension of $A$.
As in this paper we do not consider $SL_{n}$-extensions with $n\geq 4$, in all that follows it will be understood that “extendable” means $SL_{3}$-extendable. Each extendable matrix is unimodular. Theorem 4.3 contains three other equivalent characterizations of simply extendable matrices; e.g., a unimodular $2\times 2$ matrix admits diagonal reduction iff it is simply extendable.
The problem of deciding if $A\in\mathbb{M}_{2}(R)$ is simply extendable relates to classical studies of finitely generated stable free modules that aim to complete matrices in $\mathbb{M}_{n\times(n+m)}(R)$ whose $n\times n$ minors generate $R$ (e.g., see [11]), but it fits within the general problem of finding square matrices with prescribed entries and coefficients of characteristic polynomials. Concretely, if $A\in Um(\mathbb{M}_{2}(R))$ is simply extendable, its simple extensions $A^{+}$ have 5 prescribed entries (out of 9) and 2 prescribed coefficients (out of 3) of $\chi_{A^{+}}(\lambda)=\lambda^{3}-\textup{Tr}(A)\lambda^{2}+\nu_{A^{+}}\lambda-1$; the nonempty subsets
$\nu(A):=\{\nu_{A^{+}}|A^{+}\;is\;a\;simple\;extension\;of\;A\}\subseteq R$
are sampled in Remark 4.4(3) and (4) Example 7.3(1). Such general problems over fields have a long history and often complete results (cf. [4] and [7]).
In general $\Theta$ is not surjective (see Theorem 1.7(4) and Example 5.3).
If $\det(A)=0$, then $A$ is extendable iff it is simply extendable. More generally, a matrix $A\in\mathbb{M}_{2}(R)$ is extendable iff its reduction modulo $R\det(A)$ is simply extendable (see Lemma 4.1(1)). Based on this, Definition 1.1 leads to the study of 3 classes of rings to be named, in the spirit of [5] and [12], using indexed letters.
Definition 1.2.
We say that $R$ is:
(1) a ${\Pi}_{2}$ ring, if each $A\in Um(\mathbb{M}_{2}(R))$ with $\det(A)=0$ is extendable;
(2) an ${E}_{2}$ ring, if each matrix in $Um(\mathbb{M}_{2}(R))$ is extendable (i.e., $\Theta$ is surjective);
(3) an ${SE}_{2}$ ring, if each matrix in $Um(\mathbb{M}_{2}(R))$ is simply extendable.
If moreover $R$ is an integral domain, we replace ring by a domain (so we speak about $\Pi_{2}$ domains, etc.)
We recall that $R$ is called an elementary divisor ring, abbreviated as EDR, if all matrices with entries in $R$ admit diagonal reduction. Equivalently, $R$ is an EDR iff each matrix in $\mathbb{M}_{2}(R)$ is equivalent to a diagonal matrix, iff each finitely presented $R$-module is a direct sum of cyclic $R$-modules (see [10], Cor. (3.7) and Thm. (3.8); see also [20], Thm. 2.1).
Let $A\in Um(\mathbb{M}_{2}(R))$. If $A$ admits diagonal reduction, let $M,N\in{GL}_{2}(R)$ be such that $MAN=\left[\begin{array}[]{cc}1&0\\ 0&\det(A)\end{array}\right]$ and let $A^{+}:=\sigma(M^{-1})\left[\begin{array}[]{ccc}1&0&0\\ 0&\det(A)&1\\ 0&-1&0\end{array}\right]\sigma(N^{-1})$; one easily computes $\Theta(A^{+})=M^{-1}MANN^{-1}=A$ (this is a particular case of Lemma 4.1(2)), hence $A$ is simply extendable. We conclude that:
Proposition 1.3.
Each EDR is an $SE_{2}$ ring.
The following classification of ${\Pi}_{2}$ rings is proved in Section 5.
Theorem 1.4.
For a ring $R$ the following statements are equivalent:
(1) The ring $R$ is a ${\Pi}_{2}$ ring.
(2) Each matrix in $Um(\mathbb{M}_{2}(R))$ of zero determinant is non-full, i.e., the product of two matrices of sizes $2\times 1$ and $1\times 2$ (equivalently, the $R$-linear map $L_{A}$ factors as a composite $R$-linear map $R^{2}\rightarrow R\rightarrow R^{2}$).
(3) For each matrix in $Um(\mathbb{M}_{2}(R))$ of zero determinant, one (hence both) of the $R$-modules $\textup{Ker}_{A}$ and $\textup{Im}_{A}$ is isomorphic to $R$.
(4) Each projective $R$-module of rank $1$ generated by two elements is isomorphic to $R$.
For pre-Schreier domains see Section 2. Recall that an integral domain $R$ is a pre-Schreier domain iff each matrix in $\mathbb{M}_{2}(R)$ of zero determinant is non-full (see [3], Thm. 1 or [14], Lem. 1). Thus each pre-Schreier domain is a $\Pi_{2}$ domain. Similarly, if $Pic(R)$ is trivial, then $R$ is a ${\Pi}_{2}$ ring.
The following notions introduced by Bass (see [1]), Shchedryk (see [17]), and McGovern (see [13]) (respectively) will be often used in what follows.
Definition 1.5.
Let $n\in\mathbb{N}$. Recall that $R$ has:
(1) stable range n and we write $sr(R)=n$ if $n$ is the smallest natural number with the property that each $(a_{1},\ldots,a_{n},b)\in Um(R^{n+1})$ is reducible, i.e., there exists $(r_{1},\ldots,r_{n})\in R^{n}$ such that $(a_{1}+br_{1},\ldots,a_{n}+br_{n})\in Um(R^{n})$ (when there exists no $n\in\mathbb{N}$ such that $sr(R)=n$, then $sr(R):=\infty$ and the convention is $\infty>n$);
(2) (fractional) stable range 1.5 and we write $fsr(R)=1.5$ if for each $(a,b,c)\in Um(R^{3})$ with $c\neq 0$ there exists $r\in R$ such that $(a+br,c)\in Um(R^{2})$;
(3) almost stable range 1 and we write $asr(R)=1$ if for each ideal $I$ of $R$ not contained in $J(R)$, $sr(R/I)=1$.
Stable range type of conditions on all suitable unimodular tuples of a ring date back at least to Kaplansky. For instance, the essence of [8], Thm. 5.2 can be formulated in the language of this paper as follows: each triangular matrix in $Um(\mathbb{M}_{2}(R))$ is simply extendable iff for all $(a,b,c)\in Um(R^{3})$, there exists $(e,f)\in R^{2}$ such that $(ae,be+cf)\in Um(R^{2})$. The general (nontriangular) form of this reformulation is: a ring $R$ is an $SE_{2}$ ring iff for all $(a,b,c,d)\in Um(R^{4})$, there exists $(e,f)\in R^{2}$ such that $(ae+cf,be+df)\in Um(R^{2})$ (see Theorem 4.3).
We recall that if $R$ has (Krull) dimension $d$, then $sr(R)\leq d+1$ (see [1], Thm. 11.1 for the noetherian case and see [6], Cor. 2.3 for the general case). Also, if $R$ is a finitely generated algebra over a finite field of dimension $2$ or if $R$ is a polynomial algebra in 2 indeterminates over a field that is algebraic over a finite field, then $sr(R)\leq 2$ (see [19], Cors. 17.3 and 17.4).
Each ${SE}_{2}$ ring is an ${E}_{2}$ ring, but we do not know when the converse is true. However, we show that there exist $2\times 2$ matrices that are extendable but are not simply extendable (see Example 6.1). Moreover, we have (see Section 6):
Theorem 1.6.
If $sr(R)\leq 2$, then the extendable and simply extendable properties on a matrix in $\mathbb{M}_{2}(R)$ are equivalent (hence $R$ is an $SE_{2}$ ring iff it is an $E_{2}$ ring).
Based on Theorem 1.4, in Section 6 we prove the following theorem.
Theorem 1.7.
Let $R$ be an integral domain of dimension 1. Then the following properties hold:
(1) Each matrix in $Um(\mathbb{M}_{2}(R))$ with nonzero determinant is simply extendable.
(2) Each triangular matrix in $Um(\mathbb{M}_{2}(R))$ is simply extendable.
(3) The ring $R$ is a ${\Pi}_{2}$ domain iff it is an ${SE}_{2}$ (or an $E_{2}$) domain and iff $Pic(R)$ is trivial.
(4) Assume $R$ is a Dedekind domain. The ring $R$ is a ${\Pi}_{2}$ domain iff it is a principal ideal domain (PID).
Recall that $R$ is a Hermite ring in the sense of Kaplansky, if $RUm(R^{2})=R^{2}$, equivalently if each $1\times 2$ matrix with entries in $R$ admits diagonal reduction. If $R$ is a Hermite ring, then $sr(R)\in\{1,2\}$ (see [15], Prop. 8(i); see also [21], Thm. 2.1.2). As $RUm(R^{2})=R^{2}$ implies $RUm(\mathbb{M}_{2}(R))=\mathbb{M}_{2}(R)$, it follows that a Hermite ring $R$ is an EDR iff each matrix in $Um(\mathbb{M}_{2}(R))$ admits diagonal reduction (equivalently, it is simply completable, see Theorem 4.3). From the last two sentences and Theorem 1.6 we conclude:
Corollary 1.8.
Let $R$ be a Hermite ring. Then $R$ is an EDR iff it is an $E_{2}$ ring and iff it is an $SE_{2}$ ring.
For almost stable range 1 we have the following applications (see Section 6):
Corollary 1.9.
Assume that $asr(R)=1$. Then the following properties hold:
(1) Each triangular matrix in $Um(\mathbb{M}_{2}(R))$ is simply extendable.
(2) (McGovern) If $R$ is a Hermite ring, then $R$ is an EDR.
Corollary 1.9(2) was first obtained in [13], Thm. 3.7. A second proof of Corollary 1.9(2) is presented in Remark 7.2.
Part II studies determinant liftable $2\times 2$ matrices that generalize simply extendable matrices (a matrix $A\in Um(\mathbb{M}_{2}(R))$ will be called determinant liftable if there exists $B\in Um(\mathbb{M}_{2}(R))$ congruent to $A$ modulo $R\det(A)$ and with $\det(B)$=0). Part III has applications to Bézout rings (i.e., to rings whose finitely generated ideals are principal), such as each $J_{2,1}$ domain introduced in [12] is an EDR. Part IV contains universal and stability properties, complements and open problems. Parts I to IV split the manuscript https://arxiv.org/abs/2303.08413.
2. Basic terminology and properties
In what follows we will use without extra comments the following two basic properties. For $(a,b,c)\in R^{3}$ we have $(a,bc)\in Um(R^{2})$ iff $(a,b),(a,c)\in Um(R^{2})$. If $(a,b)\in Um(R^{2})$ and $c\in R$, then $a$ divides $bc$ iff $a$ divides $c$.
A ring $R$ is called pre-Schreier, if every nonzero element $a\in R$ is primal, i.e., if $a$ divides a product $bc$ of elements of $R$, there exists $(d,e)\in R^{2}$ such that $a=de$, $d$ divides $b$ and $e$ divides $c$. Pre-Schreier domains were introduced by Zafrullah in [22]. A pre-Schreier integrally closed domain was called a Schreier domain by Cohn in [2]. Every GCD domain (in particular, every Bézout domain) is Schreier (see [2], Thm. 2.4). A product of pre-Schreier domains is a pre-Schreier ring but we do not know other examples of pre-Schreier rings.
In an integral domain, an irreducible element is primal iff it is a prime. Thus an integral domain that has irreducible elements which are not prime, such as each noetherian domain which is not a UFD, is not pre-Schreier.
The inner rank of an $m\times n$ matrix over a ring is defined as the least positive integer $r$ such that it can be expressed as the product of an $m\times r$ matrix and an $r\times n$ matrix; over fields, this notion coincides with the usual notion of rank. A square matrix is called full if its inner rank equals its order, and non-full otherwise. A $2\times 2$ matrix is non-full iff its inner rank is $1$, (i.e., it has a column-row decomposition).
We consider the subsets of $Um(R^{3})$:
$T_{3}(R):=Um(R^{2})\times(R\setminus\{0\})\;\text{{and}}\;J_{3}(R):=Um(R^{2})\times(R\setminus J(R)).$
Proposition 2.1.
We have $fsr(R)=1.5$ iff for all $(a,b,c)\in T_{3}(R)$ there exists $r\in R$ such that $(a+br,c)\in Um(R^{2})$.
Proof.
The ‘only if’ part is clear. To check the ‘if’ part, let $(a,b,c)\in Um(R^{3})$ with $c\neq 0$ and let $x,y,z\in R$ be such that $ax+by+cz=1$. Thus $(a,by+cz,c)\in T_{3}(R)$ and hence there exists $r\in R$ such that $(a+byr+czr,c)\in Um(R^{2})$. This implies $(a+byr,c)\in Um(R^{2})$, thus $fsr(R)=1.5$. ∎
Proposition 2.2.
We have $asr(R)=1$ iff for all $(a,b,c)\in J_{3}(R)$ there exists $r\in R$ such that $(a+br,c)\in Um(R^{2})$.
Proof.
See [13], Thm. 3.6 for the ‘only if’ part. For the ‘if’ part, for $I$ an ideal of $R$ not contained in $J(R)$ we check that $sr(R/I)=1$. If $(a,b)\in R^{2}$ is such that $(a+I,b+I)\in Um\bigl{(}(R/I)^{2}\bigr{)}$, let $(d,e)\in R^{2}$ and $c\in I$ be such that $ad+be+c=1$. If $c\notin J(R)$, then for $(f,g):=(c,c)\in\bigl{(}I\setminus J(R)\bigr{)}\times I$ we have $(a,be+g,f)\in J_{3}(R)$. If $c\in J(R)$, then $ad+be=1-c\in U(R)$, so for $(f,g)\in\bigl{(}I\setminus J(R)\bigr{)}\times\{0\}$ we have $(a,be+g,f)\in J_{3}(R)$. If $r\in R$ is such that $(a+(be+g)r,f)\in Um(R^{2})$, then $a+I+(b+I)(er+I)\in U(R/I)$; so $sr(R/I)=1$. ∎
Corollary 2.3.
(1) If $sr(R)=1$, then $fsr(R)=1.5$.
(2) If $fsr(R)=1.5$, then $asr(R)=1$.
(3) If $asr(R)=1$, then $sr(R)\leq 2$.
Proof.
If $sr(R)=1$, then for each $(a,b,c)\in T_{3}(R)$, there exists $r\in R$ such that $a+rb\in U(R)$, so $(a+rb,c)\in Um(R^{2})$, hence $fsr(R)=1.5$ by Proposition 2.1. So part (1) holds. As $J_{3}(R)\subseteq T_{3}(R)$, part (2) follows from Propositions 2.1 and 2.2.
To check part (2), it suffices to show that each $(a,b,c)\in Um(R^{3})$ is reducible. If $b\notin J(R)$, then $R/Rb$ has stable range $1$ and $(a+Rb,c+Rb)\in Um\bigl{(}(R/Rb)^{2}\bigr{)}$; hence there exists $r\in R$ such that $a+cr+Rb\in U(R/Rb)$ and thus for $(r_{1},r_{2}):=(r,0)$ we have $(a+cr_{1},b+cr_{2})\in Um(R^{2})$. If $b\in J(R)$, then $(a,b,c)\in Um(R^{3})$ implies that $(a,c),(a,b+c)\in Um(R^{2})$ and thus for $(r_{1},r_{2}):=(0,1)$ we have $(a+cr_{1},b+cr_{2})\in Um(R^{2})$. We conclude that $(a,b,c)$ is reducible. ∎
Corollary 2.3(3) was first obtained in [13], Thm. 3.6.
We have the following ‘classical’ units and unimodular interpretations.
Proposition 2.4.
(1) For $n\in\mathbb{N}$, we have $sr(R)\leq n$ iff for each $b\in R$ the reduction modulo $Rb$ map of sets $Um(R^{n})\rightarrow Um\bigl{(}(R/Rb)^{n}\bigr{)}$ is surjective.
(2) We have $fsr(R)=1$ iff for all $(b,c)\in R^{2}$ with $c\neq 0$, the homomorphism $U(R/Rc)\rightarrow U(R/(Rb+Rc))$ is surjective.
(3) We have $asr(R)=1$ iff for all $(b,c)\in R^{2}$ with $c\notin J(R)$, the homomorphism $U(R/Rc)\rightarrow U(R/(Rb+Rc))$ is surjective.
Proof.
Parts (1) and (2) follow from definitions. To check the ‘if’ part of (3), let $(a,b,c)\in J_{3}(R)$. Then $a+Rb+Rc$ is a unit of $R/(Rb+Rc)$ and thus is the image of a unit of $R/Rc$, which is of the form $a+br+Rc$ with $r\in R$. Hence $(a+br,c)\in Um(R^{2})$. From this and Proposition 2.2 it follows that $asr(R)=1$. To check the ‘only if’ part of (3), let $a+Rb+Rc\in U(R/(Rb+Rc))$. Thus $(a,b,c)\in Um(R^{3})$. As $Rc\not\subseteq J(R)$ and we assume $asr(R)=1$, it follows that $sr(R/Rc)=1$, hence there exists $r\in R$ such that $a+br+Rc\in U(R/Rc)$ maps to $a+Rb+Rc\in U(R/(Rb+Rc))$. Thus the homomorphism $U(R/Rc)\rightarrow U(R/(Rb+Rc))$ is surjective. ∎
Corollary 2.5.
Assume that $sr(R)\leq 4$. If $R$ is an $E_{2}$ (resp. $SE_{2}$) ring, then $R/Ra$ is an $E_{2}$ (resp. $SE_{2}$) ring for all $a\in R$.
Proof.
Let $\bar{A}\in Um(\mathbb{M}_{2}(R/Ra))$. As $sr(R)\leq 4$, there exists $A\in Um(\mathbb{M}_{2}(R))$ whose reduction modulo $Ra$ is $\bar{A}$ by Proposition 2.4(1) applied to $n=4$. Let $A^{+}$ be an extension (resp. a simple extension) of $A$; its reduction modulo $Ra$ is an extension (resp. a simple extension) of $\bar{A}$. So $R/Ra$ is an $E_{2}$ (resp. $SE_{2}$) ring.∎
Example 2.6.
Let $(a,b,c)\in Um(R^{3})$ with $(b,c)\in Um(R^{2})$. Writing $a-1=be+cf$ with $e,f\in R$, for $r:=-e$ we have $(a+br,c)\in Um(R^{2})$.
Example 2.7.
Suppose $R$ is a semilocal ring. For each $b\in R$, the homomorphism $U(R)\rightarrow U(R/Rb)$ is surjective, hence $sr(R)=1$ by Proposition 2.4(1).
Example 2.8.
Let $R$ be a noetherian domain of dimension 1. For $(b,c)\in R^{2}$ with $c\neq 0$, the rings $R/Rc$ and $R/(Rb+Rc)$ are artinian, hence the homomorphism $U(R/Rc)\rightarrow U(R/(Rb+Rc))$ is surjective. From this and Proposition 2.4(2), it follows that $fsr(R)=1.5$; thus $sr(R)\leq 2$ (see Corollary 2.3(2) and (3)).
An argument similar to the one of Example 2.8 shows that each Bézout domain which is a filtered union of Dedekind domains has stable range 1.5.
Example 2.9.
We have $sr(\mathbb{Z})=2$ and $fsr(\mathbb{Z})=1.5$ and the ring $\mathbb{Z}[x]/(x^{2})$ has almost stable range 1 but does not have stable range 1.5.
3. Projective modules
For a projective $R$-module $P$ of rank $1$, let $[P]\in{Pic}(R)$ be its class in the Picard group. Let ${Pic}_{2}(R)$ be the subgroup of ${Pic}(R)$ generated by classes $[P]$ with $P$ generated by $2$ elements. Let ${Pic}_{2}(R)[2]$ be the subgroup of ${Pic}_{2}(R)$ generated by classes $[P]$ with $P\oplus P\cong R^{2}$ (so $2[P]=[R]$ and $P$ is generated by $2$ elements).
For $(\alpha_{1},\ldots,\alpha_{n})\in R^{n}$, let ${Diag}(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{M}_{n}(R)$ be the diagonal matrix whose $(i,i)$ entry is $\alpha_{i}$ for all $i\in\{1,\ldots,n\}$.
If $n\geq 2$ we do not have unit interpretations of the stable range $n$ similar to Proposition 2.4 but this is replaced by standard projective modules considerations recalled here in the form required in the sequel. If $P_{1}$ and $P_{2}$ are two projective $R$-modules of rank $1$ such that $P_{1}\oplus P_{2}\cong R^{2}$, then by taking determinants it follows that $P_{1}\otimes_{R}P_{2}\cong R$, hence $[P_{1}]=-[P_{2}]\in Pic(R)$; hence $P_{1}\cong R$ iff $P_{2}\cong R$.
[…]
We consider $R$-linear maps $g:R^{n}\rightarrow R$ such that $\pi\circ g=f$. Then $(a_{1},\ldots,a_{n},b)$ is reducible iff we can choose $g$ to be surjective. Thus, if $sr(R)\leq n$, then we can choose $g$ to be surjective.
If $A\in\mathbb{M}_{n}(R)$ is equivalent to ${Diag}(1,1,\ldots,1,b)$ and $\textup{Im}_{A}$ is the $R$-submodule $Q$ of $R^{n}$, then we can choose $g$ to be surjective and hence $(a_{1},\ldots,a_{n},b)$ is reducible.
[…]
Let $A\in Um(\mathbb{M}_{2}(R))$. Let $\bar{R}:=R/R\det(A)$ and let $\bar{A}$ be the reduction of $A$ modulo $R\det(A)$. As $\textup{Coker}_{A}$ is annihilated by $\det(A)$, we can view it as an $\bar{R}$-module isomorphic to $\textup{Coker}_{\bar{A}}$. Locally in the Zariski topology of the spectrum of $\bar{R}$, one of the entries of $\bar{A}$ is a unit and hence the matrices $\bar{A}$ and ${Diag}(1,0)$ are equivalent. Thus $\textup{Im}_{\bar{A}}$ and $\textup{Coker}_{\bar{A}}$ are projective $\bar{R}$-modules of rank $1$ generated by two elements and we have two (split) short exact sequences of projective $\bar{R}$-modules $0\rightarrow\textup{Im}_{\bar{A}}\rightarrow\bar{R}^{2}\rightarrow\textup{Coker}_{\bar{A}}\rightarrow 0$ and $0\rightarrow\textup{Ker}_{\bar{A}}\rightarrow\bar{R}^{2}\rightarrow\textup{Im}_{\bar{A}}\rightarrow 0$. From the existence of $\bar{R}$-linear isomorphisms $\bar{R}^{2}\cong\textup{Ker}_{\bar{A}}\oplus\textup{Im}_{\bar{A}}\cong\textup{Ker}_{\bar{A}}\oplus\textup{Coker}_{\bar{A}}$ it follows that $\textup{Ker}_{\bar{A}}\cong\textup{Coker}_{\bar{A}}$ are the dual of $\textup{Im}_{\bar{A}}$. Thus $\textup{Ker}_{\bar{A}}\cong\bar{R}$ iff $\textup{Im}_{\bar{A}}\cong\bar{R}$ (equivalently, $\textup{Coker}_{\bar{A}}\cong\bar{R}$), hence $[\textup{Ker}_{\bar{A}}]=[\textup{Coker}_{\bar{A}}]=-[\textup{Im}_{\bar{A}}]\in{Pic}_{2}(R)$.
4. Criteria on extending $2\times 2$ matrices
Lemma 4.1.
The following hold:
(1) A matrix $A\in\mathbb{M}_{2}(R)$ is extendable iff its reduction modulo $R\det(A)$ is simply extendable. Thus, if $\det(A)=0$, then $A$ is extendable iff it is simply extendable.
(2) For $M,N\in{GL}_{2}(R)$ and $Q\in{SL}_{3}(R)$ we have an identity
$\Theta(\sigma(M)Q\sigma(N))=M\Theta(Q)N$
and the $(3,3)$ entries of $Q$ and $\sigma(M)Q\sigma(N)$ are equivalent (thus one such entry is 00 iff the other entry is 00). Also, $\Theta(Q^{T})=\Theta(Q)^{T}$.
(3) The fact that a matrix $A\in\mathbb{M}_{2}(R)$ is (simply) extendable depends only on the double coset $[A]\in{GL}_{2}(R)\backslash\mathbb{M}_{2}(R)/GL_{2}(R)$. Moreover, $A$ is (simply) extendable iff so is $A^{T}$. So ${Im}(\Theta)$ is stable under transposition and equivalence.
Proof.
To check the ‘only if’ part of part (1), let $A\in\mathbb{M}_{2}(R)$ be extendable, with $A^{+}\in SL_{3}(R)$ an extension of it. If $A^{+}_{0}$ is obtained from $A^{+}$ by replacing the $(3,3)$ entry with 00, then the reductions of $A^{+}$ and $A^{+}_{0}$ modulo $R\det(A)$ have the same determinant $1$, and it follows that $A$ modulo $R\det(A)$ is simply extendable. To check the ‘if’ part of (1), let $B\in\mathbb{M}_{3}(R)$ be such that $\Theta(B)=A$ and its reduction modulo $R\det(A)$ is a simple extension of the reduction of $A$ modulo $R\det(A)$. Let $w\in R$ be such that $\det(B)=1+w\det(A)$. If $A^{+}\in\mathbb{M}_{3}(R)$ is obtained from $B$ by subtracting $w$ from its $(3,3)$ entry, then $A^{+}$ is an extension of $A$ as $\det(A^{+})=\det(B)-w\det(A)=1$. Thus part (1) holds. Part (2) is a simple computation, while part (3) follows directly from part (2).∎
Corollary 4.2.
If $R/Ra$ is a $\Pi_{2}$ ring for all $a\in R$, then $R$ is an $E_{2}$ ring.
Proof.
Let $A\in Um(\mathbb{M}_{2}(R))$. Its reduction modulo $R\det(A)$ has zero determinant and hence it is simply extendable as $R/R\det(A)$ is a $\Pi_{2}$ ring. From Lemma 4.1(1) it follows that $A$ is extendable. Thus $R$ is an $E_{2}$ ring.∎
Theorem 4.3.
For $A=\left[\begin{array}[]{cc}a&b\\ c&d\end{array}\right]\in Um(\mathbb{M}_{2}(R))$ the following statements are equivalent:
(1) The matrix $A$ is equivalent to the diagonal matrix ${Diag}(1,\det(A))$.
(2) The matrix $A$ is simply extendable.
(3) There exists $(e,f)\in R^{2}$ such that $(ae+cf,be+df)\in Um(R^{2})$ (note that $(e,f)\in Um(R^{2})$).
(4) There exists $(x,y,z,w)\in R^{4}$ such that $ax+by+cz+dw=1$ and the matrix $\left[\begin{array}[]{cc}x&y\\ z&w\end{array}\right]$ is non-full.
Proof.
For $(e,f,s,t)\in R^{4}$ we have two identities for the determinant
(I) $\det\left[\begin{array}[]{ccc}a&b&f\\ c&d&-e\\ -t&s&0\end{array}\right]=(be+df)t+(ae+cf)s=a(es)+b(et)+c(fs)+d(ft).$
The equivalence $(2)\Leftrightarrow(3)$ follows from the first identity. The implication $(1)\Rightarrow(2)$ was checked before Proposition 1.3.
To show that $(3)\Rightarrow(1)\wedge(4)$, let $(e,f)\in R^{2}$ be such that there exists $(s,t)\in Um(R^{2})$ with $(ae+cf)s+(be+df)t=1$. Then $(e,f),(s,t)\in Um(R^{2})$ and thus there exists $M\in{SL}_{2}(R)$ whose first row is $[e\;f]$ and there exists $N\in{SL}_{2}(R)$ whose first column is $[s\;t]^{T}$. The matrix $MAN$ has determinant $\det(A)$ and its $(1,1)$ entry is $1$, thus it is equivalent to the matrix ${Diag}(1,\det(A))$ and therefore statement $(1)$ holds by the transitivity of the equivalent relation. By taking $(x,y,z,w):=(es,et,fs,ft)$, then $\left[\begin{array}[]{cc}x&y\\ z&w\end{array}\right]=\left[\begin{array}[]{c}e\\ f\end{array}\right]\left[\begin{array}[]{cc}s&t\end{array}\right]$ is non-full and from the second identity we get that $ax+by+cz+dw=1$, thus statement (4) holds. If statement (4) holds, let $(e,f,s,t)\in R^{4}$ be such that $(x,y,z,w)=(es,et,fs,ft)$; as $ax+by+cz+dw=1$, the determinant is $1$ by the second identity, hence $A$ is simply extendable. ∎
Remark 4.4.
(1) Referring to Theorem 4.3, as $(2)\Leftrightarrow(3)$ and as a $2\times 2$ matrix has a (simple) extension iff its transpose has it, it follows that $A$ is simply extendable iff there exists $(e^{\prime},f^{\prime})\in R^{2}$ such that $(ae^{\prime}+bf^{\prime},ce^{\prime}+df^{\prime})\in Um(R^{2})$.
(2) If $A^{+}=\left[\begin{array}[]{ccc}a&b&f\\ c&d&-e\\ -t&s&0\end{array}\right]$ is a simple extension of $A=\left[\begin{array}[]{cc}a&b\\ c&d\end{array}\right]\in Um(\mathbb{M}_{2}(R))$, then the characteristic polynomial $\chi_{A^{+}}$ of $A^{+}$ is of the form
$x^{3}-{Tr}(A)x^{2}+\nu_{A^{+}}x-1$
(see Section 1 for ${Tr}(A)=a+d$ and $\nu_{A^{+}}=\det(A)+es+ft$). Thus the set of characteristic polynomials of simple extensions of $A$ is in bijection to the subset $\nu(A)\subseteq R$ introduced in Section 1 and we have
$\nu(A)=\{\det(A)+es+ft|(e,f,s,t)\in R^{4},\;a(es)+b(et)+c(fs)+d(ft)=1\}.$
(3) If $R=\mathbb{Z}$, $\{(8+11k,-5-7k)|k\in\mathbb{Z}\}$ is the solution set of the equation $17es+11ft=$ 1 in $es$ and $ft$, thus $\nu({Diag}(7,11))=\{80+4k|k\in\mathbb{Z}\}=4\mathbb{Z}$.
(4) If $d\in R$, then $\nu(Diag(1,d))=\{d+es+ft|(e,f,s,t)\in R^{4},\ es+dft=1\}$ is equal to $\{d+1-(d-1)ft|(f,t)\in R^{2}\}=2+R(d-1)$.
Example 4.5.
We check elementarily that each Bézout domain is a $\Pi_{2}$ domain. Let $A\in Um(\mathbb{M}_{2}(R))$ with $\det(A)=0$. As $A$ is equivalent to an upper triangular matrix, based on Lemma 4.1(3) we can assume that $A=\left[\begin{array}[]{cc}a&b\\ 0&d\end{array}\right]$ is upper triangular with $ad=0$. Let $p,q,r\in R$ be such that $1=ap+bq+dr$. We have $(1-ap)(1-dr)=bq$ and hence there exists $(e,s,t,v)\in R^{4}$ such that we can write $b=tv$, $1-ap=st$, $1-dr=ev$. Then $a=a(1-dr)=aev$ and $be=evt=t-dtr$, hence $be+dtr=t$. Thus the ideal $Rae+R(be+dtr)$ contains $t$ and $a$, hence it is $R$ as $1=ap+st$. From this and Theorem 4.3 (applied to $f:=tr$) it follows that $A$ is simply extendable.
Corollary 4.6.
(1) The ring $R$ is an ${SE}_{2}$ ring iff for each $(a,b,c,d)\in Um(R^{4})$ statement (3) (or (4)) of Theorem 4.3 holds.
(2) Each semilocal ring is an ${SE}_{2}$ ring.
Proof.
Part (1) follows from definitions and Theorem 4.3. Part (2) follows from part (1) and the fact that, as $R$ is semilocal, for each $(a,b,c,d)\in Um(R^{4})$ there exists $(e,f)\in R^{2}$ such that the intersection $U(R)\cap\{ae+cf,be+df\}$ is nonempty. ∎
Corollary 4.7.
Let $(a,b,c,d)\in Um(R^{4})$. The following are equivalent:
(1) The matrix $\left[\begin{array}[]{cc}a&b\\ c&d\end{array}\right]$ is extendable.
(2) There exists $(e,f)\in R^{2}$ such that $(ae+cf,be+df,ad-bc)\in Um(R^{3})$.
Similarly, (1) with simply extendable is equivalent to (2) with $(e,f)\in Um(R^{2})$.
Proof.
The ’extendable’ case and the implication $(1)\Rightarrow(2)$ in the ’simply extendable’ case follow from Lemma 4.1(1) and Theorem 4.3. We check that if there exists $(e,f)\in Um(R^{2})$ such that $(ae+cf,be+df,ad-bc)\in Um(R^{3})$, then $A$ is simply extendable. Based on Theorem 4.3 it suffices to show that we have $(ae+cf,be+df)\in Um(R^{2})$. Let $I:=R(ae+cf)+R(be+df)$ and $\mathfrak{m}$ a maximal ideal of $R$. If $ad-bc\in\mathfrak{m}$, then $(ae+cf,be+df,ad-bc)\in Um(R^{3})$ implies that $I\not\subseteq\mathfrak{m}$. If $ad-bc\notin\mathfrak{m}$, then $A$ modulo $\mathfrak{m}$ is invertible, hence $I\subseteq\mathfrak{m}$ iff $Re+Rf\subseteq\mathfrak{m}$; from this and $(e,f)\in Um(R^{2})$ we infer that $I\not\subseteq\mathfrak{m}$. As $I$ is not contained in any maximal ideal of $R$, it follows that $(ae+cf,be+df)\in Um(R^{2})$. ∎
Corollary 4.8.
The following properties hold:
(1) The ring $R$ is an ${E}_{2}$ ring iff for each $(a,b,c,d)\in Um(R^{4})$ there exists $(e,f)\in R^{2}$ such that $(ae+cf,be+df,ad-bc)\in Um(R^{3})$.
(2) The ring $R$ is an ${SE}_{2}$ ring iff for each $(a,b,c,d)\in Um(R^{4})$ there exists $(e,f)\in Um(R^{2})$ such that $(ae+cf,be+df,ad-bc)\in Um(R^{3})$.
Proof.
Both parts follow directly from definitions and the corresponding two sentences of Corollary 4.7. ∎
Example 4.9.
Let $A=\left[\begin{array}[]{cc}a&b\\ c&d\end{array}\right]\in Um(\mathbb{M}_{2}(R))$. In many simple cases one can easily prescribe $(e,f)\in R^{2}$ such that $(ae+cf,be+df)\in Um(R^{2})$ and hence conclude that $A$ is simply extendable. We include four such cases as follows.
(1) If $\{a,b,c,d\}\cap U(R)\neq\emptyset$, then we can take $(e,f)\in R^{2}$ such that $\{e,f\}=\{0,1\}$. E.g., if $a\in U(R)$, then $\left[\begin{array}[]{ccc}a&b&0\\ c&d&-1\\ 0&a^{-1}&0\end{array}\right]$ is a simple extension of $A$.
(2) If $\{(a,b),(a,c),(b,d),(c,d)\}\cap Um(R^{2})\neq\emptyset$, then we can take $(e,f,e^{\prime},f^{\prime})\in R^{4}$ such that $1\in\{ae+cf,be+df,ae^{\prime}+bf^{\prime},ce^{\prime}+df^{\prime}\}$. E.g., if $(a,b)\in Um(R^{2})$ and $s,t\in R$ are such that $as+bt=1$, then $\left[\begin{array}[]{ccc}a&b&0\\ c&d&-1\\ -t&s&0\end{array}\right]$ is a simple extension of $A$.
(3) If at least two of the entries $a$, $b$, $c$ and $d$ are in $J(R)$ (e.g., they are 00), then we can take $e,f$ such that $1\in\{ae+bf,ce+df,ae+bf+ce+df\}$. E.g., if $c=d=0$ and $s,t\in R$ are such that $as+bt=1$, then $\left[\begin{array}[]{ccc}a&b&0\\ 0&0&-1\\ -t&s&0\end{array}\right]$ is a simple extension of $A$. Similarly, if $b=c=0$ and $e,f\in R$ are such that $ae+df=1$, then $\left[\begin{array}[]{ccc}a&0&f\\ 0&d&-e\\ -1&1&0\end{array}\right]$ is a simple extension of $A$.
(4) If $a,b,c,d,f,q\in R$ are such that $aq+df=1$, then a simple extension of $\left[\begin{array}[]{cc}a&ab\\ ac&d\end{array}\right]$ is $\left[\begin{array}[]{ccc}a&ab&f\\ ac&d&-q+cf(1-b)\\ -1&1-b&0\end{array}\right]$.
In three of these examples, the $(2,3)$ entry of the simple extensions is $-1$, i.e., we can choose $e=1$. Such extensions relate to stable ranges 1 and 1.5 as follows.
Corollary 4.10.
The following properties hold:
(1) We have $sr(R)=1$ iff each upper triangular matrix $A\in Um(\mathbb{M}_{2}(R))$ has a simple extension whose $(2,3)$ entry is $-1$.
(2) We have $fsr(R)=1.5$ iff each upper triangular matrix $A\in Um(\mathbb{M}_{2}(R))$ with nonzero $(1,1)$ entry has a simple extension whose $(2,3)$ entry is $-1$.
(3) We have $asr(R)=1$ iff each upper triangular matrix $A\in Um(\mathbb{M}_{2}(R))$ with $(1,1)$ entry not in $J(R)$ has a simple extension whose $(2,3)$ entry is $-1$.
Proof.
Let $A=\left[\begin{array}[]{cc}a&b\\ 0&d\end{array}\right]\in Um(\mathbb{M}_{2}(R))$. If $a=0$, then $(b,d)\in Um(R^{2})$ and from Equation (I) it follows that $A$ has a simple extension with the $(2,3)$ entry $-1$ iff there exists $(f,t)\in R^{2}$ such that $bt+dft=1$ and hence iff there exists $f\in R$ such that $b+df\in U(R)$. Thus all these matrices $A$ with $a=0$ have a simple extension with the $(2,3)$ entry $-1$ iff $sr(R)=1$.
Similarly, if $a\neq 0$ (resp. $a\notin J(R)$), then from Equation (I) it follows that $A$ has a simple extension with the $(2,3)$ entry $-1$ iff there exists $(e,f,t)\in R^{3}$ such that $ae+bt+dft=1$ and hence iff there exists $f\in R$ such that $(b+df,a)\in Um(R^{2})$. From the definition of stable range 1.5 (resp. almost stable range 1) applied to $(b,c,a)\in Um(R^{3})$ it follows that all these matrices with $a\neq 0$ (resp. $a\notin J(R)$) have a simple extension with the $(2,3)$ entry $-1$ iff $fsr(R)=1.5$ (resp. $asr(R)=1$). ∎
5. Proof of Theorem 1.4
Proposition 5.1.
Let $A\in Um(\mathbb{M}_{2}(R))$. Then the following properties hold:
(1) If $\det(A)=0$, then $A$ is simply extendable iff it is non-full.
(2) The matrix $A$ is extendable iff its reduction modulo $R\det(A)$ is non-full.
Proof.
To prove the ‘if’ part of (1), we write $A=\left[\begin{array}[]{c}l\\ m\end{array}\right]\left[\begin{array}[]{cc}n&q\end{array}\right]$. As $A\in Um(\mathbb{M}_{2}(R))$, it follows that $(l,m),(n,q)\in Um(R^{2})$; let $(e,f),(s,t)\in R^{2}$ be such that $el+fm=sn+tq=1$. Thus $\left[\begin{array}[]{cc}e&f\end{array}\right]A=\left[\begin{array}[]{cc}n&q\end{array}\right]$. Hence $A$ is simply extendable by Theorem 4.3, a simple extension of it being $\left[\begin{array}[]{ccc}ln&lq&f\\ mn&mq&-e\\ -t&s&0\end{array}\right]$.
To prove the ‘only if’ part of (1), we assume that $\det(A)=0$ and $A$ is simply extendable. Let $(e,f)\in Um(R^{2})$ be such that $(ae+cf,be+df)\in Um(R^{2})$ (see Theorem 4.3), hence there exist pairs $(p,r)$ and $(s,t)\in Um(R^{2})$ such that $ep+fr=(ae+cf)s+(be+df)t=1$. The matrices $M:=\left[\begin{array}[]{cc}e&f\\ -r&p\end{array}\right]$ and $N:=\left[\begin{array}[]{cc}s&-be-df\\ t&ae+cf\end{array}\right]$ have determinant $1$ and a simple computation shows that, as $\det(A)=0$, we have $MAN=\left[\begin{array}[]{cc}1&0\\ w&0\end{array}\right]=\left[\begin{array}[]{c}1\\ w\end{array}\right]\left[\begin{array}[]{cc}1&0\end{array}\right]$ for some $w\in R$. Hence $A=(M^{-1}\left[\begin{array}[]{c}1\\ w\end{array}\right])(\left[\begin{array}[]{cc}1&0\end{array}\right]N^{-1})$ is non-full. Thus part (1) holds.
Part (2) follows from part (1) and Lemma 4.1(1). ∎
We are now ready to prove Theorem 1.4. Recall from Section 3 that if $A\in Um(\mathbb{M}_{2}(R))$ has zero determinant, then $\textup{Ker}_{A}$ and $\textup{Im}_{A}$ are projective $R$-modules of rank $1$ dual to each other and $\textup{Ker}_{A}\oplus\textup{Im}_{A}\cong R^{2}$. Hence $(4)\Rightarrow(3)$. The implication $(3)\Rightarrow(2)$ is obvious. The equivalence $(1)\Leftrightarrow(2)$ follows from Proposition 5.1(1).
We show that $(2)\Rightarrow(4)$. Each projective $R$-module $P$ of rank $1$ generated by $2$ elements is isomorphic to $\textup{Im}_{A}$ for some idempotent $A\in\mathbb{M}_{2}(R)$ of rank $1$ and hence unimodular of zero determinant. Assuming that (2) holds, we write $A=\left[\begin{array}[]{c}l\\ m\end{array}\right]\left[\begin{array}[]{cc}n&q\end{array}\right]$ with $(l,m),(n,q)\in Um(R^{2})$, hence $L_{A}$ is the composite of a surjective $R$-linear map $R^{2}\to R$ and an injective $R$-linear map $R\to R^{2}$. Thus $P\cong\textup{Im}_{A}\cong R$, hence $(2)\Rightarrow(4)$. Thus Theorem 1.4 holds.
Example 5.2.
For $A\in Um(\mathbb{M}_{2}(R))$ with $\det(A)=0$, we consider the $R$-submodule $K:=R\left[\begin{array}[]{c}-b\\ a\end{array}\right]+R\left[\begin{array}[]{c}-d\\ c\end{array}\right]$ of $\textup{Ker}_{A}$. For a maximal ideal $\mathfrak{m}$ of $R$, $K\not\subseteq\mathfrak{m}\mathbb{M}_{2}(R)$, hence $K\not\subseteq\mathfrak{m}\textup{Ker}_{A}$. Thus $\textup{Ker}_{A}=K$. For $(e,f)\in R^{2}$, we have $(ae+cf,be+df)\in Um(R^{2})$ iff $\textup{Ker}_{A}$ is free having $e\left[\begin{array}[]{c}-b\\ a\end{array}\right]+f\left[\begin{array}[]{c}-d\\ c\end{array}\right]$ as a generator.
Example 5.3.
Let $n\in\mathbb{N}$ and let $x_{1},\ldots,x_{n}$ be indeterminates. Let $k\in\mathbb{N}$. Let $q:=4k+1$, $r:=2k+1$ and $\theta:=i\sqrt{q}\in\mathbb{C}$. We check that $\mathbb{Z}[x_{1},\ldots,x_{n}]$ is a ${\Pi}_{2}$ ring which is not an $E_{2}$ ring. As $\mathbb{Z}[x_{1},\ldots,x_{n}]$ is a UFD, it is also a Schreier domain and hence a ${\Pi}_{2}$ ring (see Section 1). Thus it suffices to show that the matrix
$A:=\left[\begin{array}[]{cc}r&1-x_{1}\\ 1+x_{1}&2\end{array}\right]\in Um(\mathbb{M}_{2}(\mathbb{Z}[x_{1},\ldots,x_{n}]))$
is not extendable. As $\det(A)=x_{1}^{2}+q$, based on Lemma 4.1(1), it suffices to show that the image $B:=\left[\begin{array}[]{cc}r&1-\theta\\ 1+\theta&2\end{array}\right]\in Um(\mathbb{M}_{2}(\mathbb{Z}[\theta]))$ of $A$, via the composite homomorphism $\mathbb{Z}[x_{1},\ldots,x_{n}]\rightarrow\mathbb{Z}[x_{1},\ldots,x_{n}]/(x_{1}^{2}+q)\rightarrow\mathbb{Z}[\theta]$ that maps $x_{1}$ to $\theta$ and $x_{2},\ldots,x_{n}$ to 00, is not simply extendable. An argument on norms shows that the element $2\in\mathbb{Z}[\theta]$ is irreducible, i.e., $2=u(2u^{-1})$ with $u\in U(\mathbb{Z}[\theta])$ are its only product decompositions. So, as $2u^{-1}$ divides neither $1-\theta$ nor $1+\theta$, $B$ is full. So $B$ is not simply extendable by Proposition 5.1(1) and the integral domain $\mathbb{Z}[\theta]$ is not a ${\Pi}_{2}$ ring. If $14k+$ 1 is square free, then $\mathbb{Z}[\theta]$ is a Dedekind domain but not a PID.
Remark 5.4.
Statements (1) to (4) of Theorem 4.3 are stable under similarity (inner automorphisms of the $R$-algebra $\mathbb{M}_{2}(R)$) but in general they are not stable under all $R$-algebra automorphisms of $\mathbb{M}_{2}(R)$. To check this, let $R$ be such that there exists an $R$-module $P$ such that $P\oplus P=R^{2}$ but $P\ncong R$; so $[P]\in{Pic}_{2}(R)[2]$. The idempotent $A$ of $\mathbb{M}_{2}(R)$ which is a projection of $R^{2}$ on the first copy of $P$ along the second copy of $P$ satisfies $\textup{Im}_{A}=P$ and $\det(A)=0$, so it is not simply extendable by Theorem 1.4 but its image under the $R$-algebra automorphism $\mathbb{M}_{2}(R)={End}_{R}(P\oplus P)\cong\mathbb{M}_{2}(R)$ defined by $End_{R}(P)\cong R$ maps $A$ to $Diag(1,0)$. E.g., if $R$ is a Dedekind domain with ${Pic}(R)\cong\mathbb{Z}/2\mathbb{Z}$ (such as $\mathbb{Z}[\sqrt{-5}]$), then we can take $P$ to be a maximal ideal of $R$ with nontrivial class in ${Pic}(R)$.
6. Proofs of Theorems 1.6 and 1.7 and Corollary 1.9
We show that if $sr(R)\leq 2$, then each extendable $A=\left[\begin{array}[]{cc}a&b\\ c&d\end{array}\right]\in Um(\mathbb{M}_{2}(R))$ is simply extendable. As $(1)\Leftrightarrow(2)$ in Corollary 4.7, let $(e^{\prime},f^{\prime})\in R^{2}$ be such that $(ae^{\prime}+cf^{\prime},be^{\prime}+df^{\prime},ad-bc)\in Um(R^{3})$. Thus $(e^{\prime},f^{\prime},ad-bc)\in Um(R^{3})$. As $sr(R)\leq 2$, there exists $(r_{1},r_{2})\in R^{2}$ such that
$(e,f):=\bigl{(}e^{\prime}+(ad-bc)r_{1},f^{\prime}+(ad-bc)r_{2}\bigr{)}\in Um(R^{2}).$
As the ideals of $R$ generated by $ae+cf,be+df,ad-bc$ and by $ae^{\prime}+cf^{\prime},be^{\prime}+df^{\prime},ad-bc$ are equal, we have $(ae+cf,be+df,ad-bc)\in Um(R^{3})$. Thus from the last part of Corollary 4.7 it follows that $A$ is simply extendable. Thus Theorem 1.6 holds.
Example 6.1.
[…]
To prove Theorem 1.7, let $R$ be an integral domain of dimension 1. To prove part (1), let $A\in Um(\mathbb{M}_{2}(R))$ with $\det(A)\neq 0$. The ring $\bar{R}:=R/\det(A)R$ has dimension 00 and hence $Pic(\bar{R})$ is trivial. From this and Theorem 1.4 it follows that $\bar{R}$ is a $\Pi_{2}$ ring and therefore the reduction $\bar{A}$ of $A$ modulo $R\det(A)$ is simply extendable. From Theorem 1.4 it follows that $\textup{Im}_{\bar{A}}\cong R/R\det(A)$. As $R$ is an integral domain, we have two projective resolutions $0\rightarrow R\det(A)\rightarrow R\rightarrow\textup{Im}_{\bar{A}}\rightarrow 0$ and $0\rightarrow\det(A)R^{2}\rightarrow\textup{Im}_{A}\rightarrow\textup{Im}_{\bar{A}}\rightarrow 0$ of $R$-modules. From this and Section 3 applied to $b=\det(A)$ (recall that $sr(R)\leq 2$), it follows that $A$ is equivalent to ${Diag}(1,\det(A))$ and hence is simply extendable (see Theorem 4.3).
Part (2) holds as triangular matrices in $Um(\mathbb{M}_{2}(R))$ have either two 00 entries or nonzero determinants and thus are simply extendable by Example 4.9(3) or part (1). The first ‘iff’ of part (3) follows from part (1) and definitions. The isomorphism classes of projective $R$-modules of rank 1 are the isomorphisms classes of nonzero ideals of $R$ which locally in the Zariski topology are principal; as for each $a\in R\setminus\{0\}$, $\dim(R/Ra)=0$ and hence $Pic(R/Ra)$ is trivial, all such nonzero ideals are generated by 2 elements. From this and Theorem 1.4 it follows that $R$ is a $\Pi_{2}$ domain iff $Pic(R)$ is trivial. Hence part (3) holds. As PIDs are precisely Dedekind domains with trivial Picard groups, part (4) follows from the second ‘iff’ of part (3). Thus Theorem 1.7 holds.
Remark 6.2.
The existence of an extension of a matrix in $Um(\mathbb{M}_{2}(R))$ does not depend only on the set of its entries. This is so as, referring to Example 5.3, the matrix $\left[\begin{array}[]{cc}1+\theta&1-\theta\\ r&2\end{array}\right]\in Um(\mathbb{M}_{2}(\mathbb{Z}[\theta]))$ has the same entries as $B$, has nonzero determinant, and it is extendable (see Theorem 1.7(1)).
To prove Corollary 1.9, we assume that $asr(R)=1$. Each matrix $\left[\begin{array}[]{cc}a&b\\ 0&c\end{array}\right]\in Um(\mathbb{M}_{2}(R))$ is simply extendable by Example 4.9(3) if $a\in J(R)$ and by Corollary 4.10(3) if $a\notin J(R)$. So part (1) holds. To prove part (2), as $R$ is also a Hermite ring, each $A\in Um(\mathbb{M}_{2}(R))$ is equivalent to a triangular matrix and hence from part (1) and Lemma 4.1(3) it follows that $A$ is simply extendable and thus admits diagonal reduction by Theorem 4.3. As $\mathbb{M}_{2}(R)=RUm(\mathbb{M}_{2}(R))$ it follows that each matrix in $\mathbb{M}_{2}(R)$ admits diagonal reduction, hence is equivalent to a diagonal matrix. Thus $R$ is an EDR. Hence Corollary 1.9 holds.
7. Explicit computations for integral domains
Let $A:=\left[\begin{array}[]{cc}a&b\\ c&d\end{array}\right]\in Um(\mathbb{M}_{2}(R))$ be such that we can write $a=ga^{\prime}$, $c=gc^{\prime}$, $b=hb^{\prime}$, $d=hd^{\prime}$ with $a^{\prime},b^{\prime},c^{\prime},d^{\prime},g,h\in R$ and $(a^{\prime},c^{\prime}),(b^{\prime},d^{\prime})\in Um(R^{2})$. We have $(g,h)\in Um(R^{2})$. Let $e^{\prime},f^{\prime}\in R$ be such that $a^{\prime}e^{\prime}+c^{\prime}f^{\prime}=1$. Let $l:=b^{\prime}c^{\prime}-a^{\prime}d^{\prime}\in R$ and $m:=b^{\prime}e^{\prime}+d^{\prime}f^{\prime}\in R$; note that $\det(A)=-ghl$. As $[\begin{array}[]{cc}l&m\end{array}]=[\begin{array}[]{cc}b^{\prime}&d^{\prime}\end{array}]\left[\begin{array}[]{cc}c^{\prime}&e^{\prime}\\ -a^{\prime}&f^{\prime}\end{array}\right]$, the matrix $\left[\begin{array}[]{cc}c^{\prime}&e^{\prime}\\ -a^{\prime}&f^{\prime}\end{array}\right]$ has determinant 1, and $({b}^{\prime}{,d}^{\prime})\in Um(R^{2})$, it follows that $(l,m)\in Um(R^{2})$.
Let $(e,f,w)\in R^{3}$ be such that $ae+cf=gw$. If $R$ is an integral domain and $g\neq 0$, there exists $v\in R$ such that $(e,f)=(we^{\prime}+c^{\prime}v,wf^{\prime}-a^{\prime}v)$, hence
$be+df=h(b^{\prime}e+d^{\prime}f)=h(b^{\prime}we^{\prime}+b^{\prime}c^{\prime}v+d^{\prime}wf^{\prime}-a^{\prime}d^{\prime}v)=hw(b^{\prime}e^{\prime}+d^{\prime}f^{\prime})+hv(b^{\prime}c^{\prime}-a^{\prime}d^{\prime})$
is equal to $h(wm+vl)$. Thus $(ae+cf,be+df)=\bigl{(}gw,h(wm+vl)\bigr{)}$.
Proposition 7.1.
Let $R$ be an integral domain. Let $A=\left[\begin{array}[]{cc}a&b\\ c&d\end{array}\right]\in Um(\mathbb{M}_{2}(R))$ be such that the above notation $g,h,a^{\prime},b^{\prime},c^{\prime},d^{\prime}$ applies and let $(e^{\prime},f^{\prime},l,m)\in R^{4}$ be obtained as above. We assume $g\neq 0$.
(1) The matrix $A$ is simply extendable iff there exists $(w,v)\in R^{2}$ such that $(g,wm+vl),(w,hvl)\in Um(R^{2})$, in which case a simple extension of $A$ is
$\left[\begin{array}[]{ccc}a&b&wf^{\prime}-a^{\prime}v\\ c&d&-we^{\prime}-c^{\prime}v\\ -t&s&0\end{array}\right]$
where $s,t\in R$ are such that $gws+h(wm+vl)t=1$.
(2) If the intersection $\{(g,l),(g,m),(h,l),(h,m)\}\cap Um(R^{2})$ is nonempty (e.g., if $hlm=0$), then $A$ is simply extendable and $w,v\in R$ are given by formulas.
Proof.
There exists $(e,f)\in R^{2}$ such that $(ae+cf,be+df)\in Um(R^{2})$ iff there exists $(w,v)\in R^{2}$ such that $\bigl{(}gw,h(wm+vl)\bigr{)}\in Um(R^{2})$ (see above). We have $\bigl{(}gw,h(wm+vl)\bigr{)}\in Um(R^{2})$ iff $\bigl{(}g,h(wm+vl)\bigr{)},\bigl{(}w,h(wm+vl)\bigr{)}\in Um(R^{2})$. As $(g,h)\in Um(R^{2})$, we have $\bigl{(}g,h(wm+vl)\bigr{)}\in Um(R^{2})$ iff $(g,wm+vl)\in Um(R^{2})$; moreover, $Rw+Rh(wm+vl)=Rw+Rhvl$. Thus $\bigl{(}gw,h(wm+vl)\bigr{)}\in Um(R^{2})$ iff $(g,wm+vl),(w,hvl)\in Um(R^{2})$. Based on the ‘iff’ statements of this paragraph and Theorem 4.3, it follows that part (1) holds.
To check part (2), we first notice that if $hlm=0$, then the intersection is nonempty; e.g., if $h=0$, then $g\in U(R)$ and hence $(g,l),(g,m)\in Um(R^{2})$. Based on part (1) it suffices to show that in all four possible cases, we can choose $(w,v)\in R^{2}$ such that $(g,wm+vl),(w,hvl)\in Um(R^{2})$.
If $(g,l)\in Um(R^{2})$, for $(w,v):=(g,1)$ we have $Rg+R(wm+vl)=Rg+Rl=R$. As $(g,h),(g,l)\in Um(R^{2})$, also $(w,hvl)=(g,hl)\in Um(R^{2})$.
If $(g,m)\in Um(R^{2})$, for $(w,v):=(1,0)$ we have $(g,wm+vl)=(g,m)\in Um(R^{2})$ and $(w,hvl)=(1,0)\in Um(R^{2})$.
If $(h,m)\in Um(R^{2})$, then $(hl,m)\in Um(R^{2})$ and there exists $(w,v^{\prime})\in R^{2}$ such that $wm+hv^{\prime}l=1$; so $Rw+Rhv^{\prime}l=R$. For $v:=hv^{\prime}$ we have $wm+vl=1$ and so $(g,wm+vl)\in Um(R^{2})$ and $(w,hvl)=(w,h^{2}v^{\prime}l)\in Um(R^{2})$ as $(w,hv^{\prime}l)\in Um(R^{2})$.
If $(h,l)\in Um(R^{2})$, let $(p,q)\in R^{2}$ be such that $1=lp+mq$. For $w:=hq+l$ and $v:=hp-m$ we compute $wm+vl=h(lp+mq)+ml-ml=h$, so $(g,wm+vl)=(g,h)\in Um(R^{2})$. Also, $Rw+Rh=R(hq+l)+Rh=Rl+Rh=R$ and $Rw+Rvl$ contains $wm+vl=h$ and hence it contains $Rw+Rh=R$. Thus $(w,vl)\in Um(R^{2})$. As $(w,h),(w,vl)\in Um(R^{2})$ it follows that $(w,hvl)\in Um(R^{2})$. ∎
Remark 7.2.
We include a second proof of Corollary 1.9(2) for Bézout domains. Assume $R$ is a Bézout domain with $asr(R)=1$. Based on the equivalence of statements (1) and (2) of Theorem 4.3, it suffices to show that each matrix $A=\left[\begin{array}[]{cc}a&b\\ c&d\end{array}\right]\in Um(\mathbb{M}_{2}(R))$ is simply extendable. As $R$ is a Hermite domain, the notation of this section applies. As $(g,h)\in Um(R^{2})$, by the symmetry between the pairs $(a,c)$ and $(b,d)$, we can assume that $g\notin J(R)$. As $(m,l,g)\in J_{3}(R)$ and $asr(R)=1$, there exists $v\in R$ such that $(m+lv,g)\in Um(R^{2})$. We take $w:=1$. Hence $(w,hvl)\in Um(R^{2})$ and $(g,wm+lv)=(g,m+lv)\in Um(R^{2})$. From Proposition 7.1(1) it follows that $A$ is simply extendable.
Example 7.3.
For $A=\left[\begin{array}[]{cc}a&b\\ c&d\end{array}\right]\in Um(\mathbb{M}_{2}(\mathbb{Z}))$, its simple extensions $A^{+}_{(e,f,s,t)}=\left[\begin{array}[]{ccc}a&b&f\\ c&d&-e\\ -t&s&0\end{array}\right]$ are parameterized by the set (see Equation (I))
$\gamma_{A}:=\{(e,f,s,t)\in\mathbb{Z}^{4}|a(es)+b(et)+c(fs)+d(ft)=1\}.$
The below examples were (initially) exemplified using a code written for $R=\mathbb{Z}$ by the second author. We have $\nu_{A^{+}_{(e,f,s,t)}}=\chi_{A^{+}_{(e,f,s,t)}}^{\prime}(0)=\det(A)+es+ft$.
(1) If $a=0$ and $d=1+b+c$, then we can take $A^{+}=\left[\begin{array}[]{ccc}0&b&-1\\ c&1+b+c&-1\\ 1&1&0\end{array}\right].$
Concretely, suppose $(b,c)=(3,2)$; then $d=6$, $\det(A)=-6$, $\gamma_{A}=\{(e,f,s,t)\in\mathbb{Z}^{4}|3et+2fs+6ft=1\}$ and $\nu(A)=\{-6+es+ft|(e,f,s,t)\in\gamma_{A}\}$.
To solve the equation $13et+2fs+6ft=$ 1, let $w:=et+2ft$. We get $12fs+3w=$ 1 with general solution $fs=-1+3k$, $w=1-2k$, where $k\in\mathbb{Z}$. The general solution of the equation $et+2ft=1-2k$ is $et=2k-1-2l$ and $ft=1-2k+l$, where $l\in\mathbb{Z}$. Let $m:=l-2k+1$. It follows that $ft=m,\;et=1-2k-2m,\;fs=-1+3k$ and the only constraint is that $ft=m$ divides $etfs=(3k-1)(1-2k-2m)$, i.e., divides $(3k-1)(2k-1)$. As $es+ft=m+\dfrac{-6k^{2}+5k-1}{m}-6k+2$, it follows that
$\nu(A)=\{-4+m-6k+\dfrac{-6k^{2}+5k-1}{m}|(m,k)\in\mathbb{Z}^{2},\;m\;{\textup{divides}}\;-6k^{2}+5k-1\}.$
(2) If $c=0$ and $d=1-a+b$, then we can take $A^{+}=\left[\begin{array}[]{ccc}a&b&-1\\ 0&1-a+b&-1\\ 1&1&0\end{array}\right]$ (cf. Corollary 4.8(2): for simple extensions of upper triangular matrices with nonzero $(1,1)$ entries over rings $R$ with $fsr(R)=1.5$, we can choose $e=1$).
Concretely, suppose $(a,b)=(6,-10)$; hence $d=15$ and $\det(A)=-90$. Thus $\gamma_{A}=\{(e,f,s,t)\in\mathbb{Z}^{4}|6es-10et-15ft=1\}$.
To solve the equation $16es-10et-15ft=$ 1, let $w:=2et+3ft$. We get $16es-5w=$ 1 with general solution $es=1+5k$, $w=1+6k$, where $k\in\mathbb{Z}$. Then $2et+3ft=1+6k$ has the general solution $et=-1-6k+3l$, $ft=1+6k-2l$, where $l\in\mathbb{Z}$. Let $m:=l-2k$. It follows that $es=1+5k,\;et=-1+3m,\;ft=1+2k-2m$ are subject to the only constraint that $et=-1+3m$ divides $esft=(1+5k)(1+2k-2m)$, i.e., it divides $(1+5k)(2k+m)$. As $es+ft=2+7k-2m$, it follows that
$\nu(A)=\{-88+7k-2m|(m,k)\in\mathbb{Z}^{2},\;-1+3m\;{\textup{divides}}\;(1+5k)(2+2k+m)\}.$
For $m=0$ (resp. $m=1$) it follows that $\nu(A)\supseteq 3+7\mathbb{Z}$ (resp. $\nu(A)\supseteq 1+14\mathbb{Z}$).
The matrices $Diag(6,-15),A_{0}:=A\in\mathbb{M}_{2}(\mathbb{Z}[\dfrac{1}{21}])$ are similar and one checks that $\nu(A_{0})=-90+\dfrac{1}{6}+\dfrac{7}{2}+2\mathbb{Z}[\dfrac{1}{21}]=-86-\dfrac{1}{3}+2\mathbb{Z}[\dfrac{1}{21}]$.
(3) If $A=$ $\left[\begin{array}[]{cc}15&6\\ 10&14\end{array}\right]$, we can take $A^{+}=\left[\begin{array}[]{ccc}15&6&-2\\ 10&14&1\\ -1&-1&0\end{array}\right]$; indeed we have $\det(A^{+})=1\cdot 15-1\cdot 6+2\cdot 10-2\cdot 14=1$. The entries of $A$ use double products of the primes $2,3,5,7$. We have $\gamma_{A}=\{(e,f,s,t)\in\mathbb{Z}^{4}|15es+6et+10fs+14ft=1\}$, $\det(A)=150$, $\nu_{A^{+}}=149$, and $\nu(A)=\{150+es+ft|(e,f,s,t)\in\gamma_{A}\}$.
To solve the equation $115es+6et+10fs+14ft=$ 1, let $x:=5es+2et$ and $y:=5fs+7ft$, so we get $13x+2y=$ 1 with general solution $x=1+2k$, $y=-1-3k$, where $k\in\mathbb{Z}$. Then $5es+2et=1+2k$ has the general solution $es=1+2k+2l$, $et=-2(1+2k)-5l$, where $l\in\mathbb{Z}$, and $5fs+7ft=-1-3k$ has the general solution $fs=3(-1-3k)+7r$, $ft=-2(-1-3k)-5r$, where $r\in\mathbb{Z}$. Let $o:=k+l$, so $et=-2+4l-4o-5l=:q$. Thus $es=1+2o$, $et=q$, $l=-2-q-4o$, $k=2+q+5o$, therefore $fs=-3-18-9q-45o+7r$ and $ft=2+12+6q+30o-5r$. Let $p:=r-6o-q-3$. Thus $ft=-5p+q-1$ and $fs=-3o-2q+7p$. As $(es)(ft)=(et)(fs)$, we have an identity $(1+2o)(q-1-5p)=q(7p-3o-2q)$, which can be rewritten as $o(-2-10p+5q)=1+5p-q-2q^{2}+7pq$ and which for $q=2p$ becomes $-2o=6p^{2}+3p+1$, requiring $p$ to be odd. For $q=2p$, $es+ft=2o+q-5p$ becomes $12o-3p=-6p^{2}-6p-$ 1. Thus
$\nu(A)=\{150+2o+q-5p|(o,q,p)\in\mathbb{Z}^{3},\;o(-2-10p+5q)=1+5p-q-2q^{2}+7pq\}$
contains the set $\{150-6p^{2}-6p-1|p-1\in 2\mathbb{Z}\}$.
(4) If $A=\left[\begin{array}[]{cc}30&42\\ 70&105\end{array}\right]$, we can take $A^{+}=\left[\begin{array}[]{ccc}30&42&1\\ 70&105&3\\ 1&1&0\end{array}\right]$; indeed we have $\det(A^{+})=-3\cdot 30+3\cdot 42+1\cdot 70-1\cdot 105=1$. The entries of $A$ use triple products of the primes $2,3,5,7$, hence they are not pairwise coprime.
Acknowledgement. The third author would like to thank SUNY Binghamton for good working conditions and Ofer Gabber for pointing out a glitch in one of the references. Competing interests: the authors declare none.